# A Robust Real-Time Lane Detection Method with Fog-Enhanced Feature Fusion for Foggy Conditions

### v2

@@ Abstract
Index Terms:
ADAS, intelligent vehicle, lane detection, real-time, foggy lane detection dataset, foggy lane detection
I Introduction
@@
Under good weather and clear lane markings, current lane detection algorithms perform well. However, in complex road environments and poor weather, such as fog, lane markings can become blurred, image quality may decrease, and global environmental information is often obscured, which makes lane detection more challenging. In these conditions, existing algorithms often struggle and may fail, leading to an increased risk of traffic accidents with more severe consequences. Ensuring safe driving in these conditions calls for the development of high-precision, highly robust lane detection methods optimized for foggy environments. As shown in Fig. 1, the in-vehicle camera captures real-time road images, which are preprocessed by edge devices and then sent to the lane detection module. Detected lane information is fused with data from other sensors, such as radar and GPS, along with real-time traffic data from the cloud. This information is processed in the vehicle’s computing unit, allowing the system to make adaptive driving decisions—such as accelerating, braking, or steering—to ensure safety in challenging environments like fog.
@@
Owing to deep learning’s superior performance in image-based tasks, it now serves as a fundamental approach for many vision tasks. In lane detection, deep learning-based approaches have become mainstream due to their strong feature extraction capabilities and generalization. Compared to traditional methods, deep learning provides higher robustness, better adapting to complex road environments and effectively handling cases with missing or occluded lane markings. LaneNet[4] approaches lane detection as a combination of semantic segmentation and post-processing, first generating a semantic segmentation map of the lanes and then using density clustering to identify each lane instance. However, because of the large number of pixels generated during segmentation and the inability of clustering to leverage GPU acceleration, its real-time performance remains limited. LaneAF[5], building on the semantic segmentation map, introduced a clustering algorithm based on horizontal and vertical vector fields to accelerate post-processing and used DLA[6] as the backbone network to integrate multi-level semantic and spatial information, enhancing feature representation. SCNN[7] leveraged the elongated structure of lane lines by replacing traditional layer-wise convolution with slice convolution, enabling cross-row and cross-column information transfer, which makes it well-suited to line-like lane structures. However, its large parameter size prevents real-time performance. Overall, current networks, whether for semantic or instance segmentation, mostly rely on pixel-by-pixel classification, though lane detection does not necessarily require dense pixel segmentation. To address this, Qin et al.[8] reframed lane detection as a row-based segmentation problem using global features, significantly reducing computational cost and improving detection speed by selecting specific rows and classifying each pixel in the row. This method also leverages the large receptive field of global features, enabling it to handle complex scenarios with missing or occluded lanes.
@@
Apart from segmentation, lane detection also includes detection-based methods, many of which borrow the Anchor-based Methods from object detection[9, 10, 11, 12, 13, 14, 15]. However, due to the elongated shape of lane lines, traditional bounding boxes are not ideal. In lane detection, anchors are often a series of preset starting points and straight lines with different slopes, represented as point sets, with the network predicting the relative distance between the lane line and anchor to optimize line shape and position. PointLaneNet[16] first represents lane lines as a series of points, taking the intersections of evenly spaced horizontal lines as coordinates $\{x_{1},x_{2},\ldots,x_{n}\}$ to describe the lane. Like other YOLO-based methods[9, 10, 17, 18, 19, 20, 21], PointLaneNet divides the image into multiple grids, predicts lane position and classification at each grid center, and removes duplicate predictions through post-processing. However, this grid-based prediction approach has limitations for elongated lanes, reducing detection accuracy. LaneATT[22] addresses this issue by noting the high correlation between lane positions and introducing global features to infer lane position, especially in cases of occlusion or missing lane markers. LaneATT[22] introduced an anchor-based attention mechanism that significantly improves detection speed and accuracy by utilizing global information. CLRNet[23] demonstrated the complementary roles of low- and high-dimensional features in lane detection, proposing Cross Layer Refinement (CLR) to dynamically adjust anchor points through network learning, removing the limitations of fixed anchors to optimize lane detection further. FLAMNet[24] proposes an adaptive line anchor framework integrated with a dual-branch network design to improve the accuracy and adaptability of capturing intricate lane geometries. By leveraging the Decomposed Self-Attention (DSA) mechanism alongside a Horizontal Feature Fusion Module (HFFM), the model enhances the perceptual range of line anchors in lane detection tasks. This approach strengthens robustness against occlusions and incomplete lane markings in challenging environments.
@@
Another approach employs domain adaptation techniques to address the style gap between foggy scenes and clear-weather images[27, 28, 29]. Hnewa et al.[30] assumed that there is domain shift between images captured in normal and adverse weather conditions, proposing a multi-scale domain-adaptive YOLO that supports domain adaptation at different levels in the feature extraction stage. Li et al.[31] proposed a novel transfer learning-based framework for fog-adaptive object detection, which integrates dual-level adaptation (image-level and object-level) to mitigate domain shifts caused by foggy conditions. Notably, the method only requires labeled clear-weather images and unlabeled foggy data for training, eliminating the dependency on annotated foggy datasets.
@@ II Related Work
II-A Segmentation-based Lane Detection
@@
II-B Curve-based Lane Detection
@@
II-C Anchor-based Lane Detection
@@
II-D Row-wise-based Lane Detection
@@
III Data Constrution
Due to the data-driven nature of deep learning models, the quality, size, and diversity of the datasets largely determine the applicability and performance of such models. In the task of lane detection, models trained solely on clear weather data struggle to adapt to low-visibility foggy conditions. Issues such as blurred lane boundaries, faded lane colors, and reduced image quality often arise in fog, making lane detection more challenging. Therefore, training with a foggy weather lane dataset can significantly improve an autonomous driving system’s ability to handle such conditions, enhancing both detection accuracy and overall reliability. However, there is currently no publicly available foggy weather lane detection dataset. To address this gap, we have collected and constructed a foggy weather lane detection dataset called FoggyLane. In addition, we generated two supplementary datasets, FoggyCULane and FoggyTusimple, by using fog modeling techniques on existing CULane [7] and Tusimple [32] datasets to test the generalizability of our methods.
@@ III Data Constrution
III-A FoggyLane
In the field of intelligent connected vehicle ADAS systems, existing lane detection algorithms have shown good performance under favorable weather conditions. However, their performance remains significantly inadequate under adverse weather conditions, such as fog. Given that fog is a high-risk scenario for traffic accidents, the limited visibility drastically reduces the driver’s reaction time and decision-making ability, making accidents more likely. However, current benchmark datasets for lane detection, including CULane[7], Tusimple[32], and LLAMAS[41], do not include data from foggy scenarios. To address this gap, we have developed a lane detection dataset, FoggyLane, specifically designed for foggy conditions. Given the randomness of foggy weather, data collection was carried out using a monocular forward-facing camera[42] at different times and locations. This includes the morning of October 26, 2023, in Suzhou, Jiangsu Province, China, and two sessions in Urumqi, Xinjiang Uygur Autonomous Region, China: the morning of December 9, 2023, and the night of January 2, 2024. In these selected representative traffic scenarios, we sampled one frame every frames, resulting in a total of 1,086 RGB images: 444 images (resolution: 1920×1080), 102 images (resolution: 2560×1920), and 540 images (resolution: 2304×1296). To simplify operations and enhance usability during training, we resized all images to a uniform resolution of 1640×590.
However, since the duration of foggy weather conditions is limited, the number of collected foggy weather images is restricted. To improve the model’s performance, we extended our dataset by incorporating additional foggy driving footage. Specifically, we downloaded two foggy driving videos from Bilibili[43], extracting road-driving segments at 20-frame intervals, obtaining 44 and 293 frames, respectively. This added a total of 337 additional RGB images to the dataset. The original resolution of these images was 1920×1080, and they were resized to 1640×590 for consistency.
@@ III-A FoggyLane
Meanwhile, We compared FoggyLane with existing lane detection datasets. As shown in Table I, the Caltech-Lanes and Tusimple datasets only cover single traffic scenarios. The LLAMAS [41] dataset focuses on highway scenarios, but its annotations are generated from high-precision maps rather than manually annotated. While the ApolloScape [44] and BDD100k [45] datasets contain multiple traffic scenarios, they lack instance-level annotations, and ApolloScape [44] does not provide publicly available test set annotations. In comparison, although CULane [7] and CurveLanes [46] include a wide range of traffic scenes with comprehensive annotations, they still lack foggy weather scenarios. For lane detection tasks, models trained only on clear weather conditions struggle to handle the challenges posed by fog, such as blurred lane boundaries, faded colors, and reduced image quality. To address these limitations, we developed FoggyLane, a comprehensive foggy weather lane detection dataset that includes 1,423 images featuring various challenging scenarios such as foggy roads during the day, curves, crossroads, and night fog. The dataset is meticulously annotated using the LabelMe tool, providing instance-level annotations, and it is available in Tusimple [32] and CULane [7] formats for broader accessibility.
TABLE I: Comparison Between Our Dataset and Existing Lane Detection Datasets
@@ III-B FoggyCULane and FoggyTusimple
IV Methods
IV-A Overall structure
@@
IV-B Global Feature Fusion Module
In recent years, Transformer models have achieved significant success in natural language processing, leading researchers to explore their applications in computer vision. With self-attention mechanisms, Transformer models can capture dependencies between different regions and pixels within an image, thus facilitating a better understanding of global image information. As a result, Transformer models have demonstrated outstanding modeling capabilities and performance improvements in various traffic vision tasks[54, 55, 56]. However, applying Transformers to the visual domain still faces some challenges. Firstly, traditional Transformer architectures lack the hierarchical structure and down-sampling operations found in CNN models, making them less adaptable to changes in resolution and receptive field, which may hinder handling of complex visual entities. Additionally, calculating global attention at high resolutions can be computationally intensive.
@@ IV-B Global Feature Fusion Module
IV-C Kernel Feature Fusion Module
@@
IV-D Low-level Edge Enhanced Module
Images in foggy conditions differ from those captured under good lighting primarily due to their blurriness and low contrast. This effect renders the edges of objects and details within the scene indistinct and hard to detect, impacting the algorithm’s ability to understand and recognize the scene accurately. Edge information, beyond simply marking object boundaries, also implicitly contains structural and spatial information about objects, forming a critical foundation for tasks such as object detection, segmentation, and tracking[62]. Using the Canny edge detection algorithm with identical threshold settings, we processed both foggy and normal images, as shown in the Fig. 10. It is evident that fog significantly degrades edge information, particularly affecting lane line edges, which become nearly indiscernible. Given that lane detection relies heavily on edge information—lane positioning is typically determined by the boundary between lane edges and the road background—both global scene information and detailed edge information are essential for accurate lane detection.
@@ IV-D Low-level Edge Enhanced Module
TABLE II: Comparison with Advanced Lane Detection on FoggyLane dataset.
@@
Specifically, we add an auxiliary path in the network to learn the edge features of the input image. First, we apply the Canny operator to the input image for edge detection, supplemented by lane line labels to reconstruct any missing lane edges, generating the training labels for the edge enhancement module. To avoid interfering with high-level semantic extraction and subsequent detection segmentation, while strengthening the low-level network’s learning of edge textures and other low-dimensional features, we introduce a branch at the second stage of the backbone network. This branch sequentially applies a 3 × 3 and a 1 × 1 convolution. We then compute the cross-entropy loss between the branch output and the edge labels. The detailed structure is illustrated in the bottom-left of the Fig. 8. Notably, this auxiliary enhancement path is only active during training, thus not increasing computation or processing time during deployment.
@@
TABLE III: Comparison with Advanced Lane Detection on FoggyCULane dataset
@@ V-A Datasets
FoggyLane is a lane detection dataset specifically designed for foggy scenarios, containing 1,423 frames of foggy lane images for training and testing. All images are of resolution 1640x590 pixels.
FoggyCULane is synthesized from the large-scale challenging lane detection dataset CULane using a foggy image modeling approach. It maintains the same number of images as CULane, which includes 88,880 images for training and 34,680 images for testing. The resolution of all images is also 1640x590 pixels.
FoggyTusimple is synthesized from one of the most widely used benchmark datasets for lane detection, Tusimple, through the foggy image modeling method. At the same time, the dataset follows the same structure as Tusimple, comprising 3,268 training images, 358 validation samples, and 2,782 test cases. All visual data maintains a consistent resolution of 1280×720 pixels.
Figure 13: Several metrics of our model’s performance[63] during the training process on the FoggyLane and FoggyTusimple datasets, with results on FoggyLane on the left and FoggyTusimple on the right.
V-B Evaluation Metric

### v3

@@ Abstract
keywords:
[…]
1 Introduction
@@
Under good weather and clear lane markings, current lane detection algorithms perform well. However, in complex road environments[1] and poor weather, such as fog, lane markings can become blurred, image quality may decrease, and global environmental information is often obscured, which makes lane detection more challenging. In these conditions, existing algorithms often struggle and may fail, leading to an increased risk of traffic accidents with more severe consequences[2]. Ensuring safe driving in these conditions calls for the development of high-precision, highly robust lane detection methods optimized for foggy environments. As shown in Fig. 1, the in-vehicle camera captures real-time road images, which are preprocessed by edge devices and then sent to the lane detection module. Detected lane information is fused with data from other sensors, such as radar and GPS, along with real-time traffic data from the cloud. This information is processed in the vehicle’s computing unit, allowing the system to make adaptive driving decisions—such as accelerating, braking, or steering—to ensure safety in challenging environments like fog.
@@
Owing to deep learning’s superior performance in image-based tasks, it now serves as a fundamental approach for many vision tasks[6, 7]. In lane detection, deep learning-based approaches[8] have become mainstream due to their strong feature extraction capabilities and generalization. Compared to traditional methods, deep learning provides higher robustness, better adapting to complex road environments and effectively handling cases with missing or occluded lane markings. LaneNet[9] approaches lane detection as a combination of semantic segmentation and post-processing, first generating a semantic segmentation map of the lanes and then using density clustering to identify each lane instance. However, because of the large number of pixels generated during segmentation and the inability of clustering to leverage GPU acceleration, its real-time performance remains limited. LaneAF[10], building on the semantic segmentation map, introduced a clustering algorithm based on horizontal and vertical vector fields to accelerate post-processing and used DLA[11] as the backbone network to integrate multi-level semantic and spatial information, enhancing feature representation. SCNN[12] leveraged the elongated structure of lane lines by replacing traditional layer-wise convolution with slice convolution, enabling cross-row and cross-column information transfer, which makes it well-suited to line-like lane structures. However, its large parameter size prevents real-time performance. Overall, current networks, whether for semantic or instance segmentation, mostly rely on pixel-by-pixel classification, though lane detection does not necessarily require dense pixel segmentation. To address this, Qin et al.[13] reframed lane detection as a row-based segmentation problem using global features, significantly reducing computational cost and improving detection speed by selecting specific rows and classifying each pixel in the row. This method also leverages the large receptive field of global features, enabling it to handle complex scenarios with missing or occluded lanes.
@@
Apart from segmentation, lane detection also includes detection-based methods, many of which borrow the Anchor-based Methods from object detection[14, 15, 16, 17]. However, due to the elongated shape of lane lines, traditional bounding boxes are not ideal. In lane detection, anchors are often a series of preset starting points and straight lines with different slopes, represented as point sets, with the network predicting the relative distance between the lane line and anchor to optimize line shape and position. PointLaneNet[18] first represents lane lines as a series of points, taking the intersections of evenly spaced horizontal lines as coordinates $\{x_{1},x_{2},\ldots,x_{n}\}$ to describe the lane. Like other YOLO-based methods[14, 15], PointLaneNet divides the image into multiple grids, predicts lane position and classification at each grid center, and removes duplicate predictions through post-processing. However, this grid-based prediction approach has limitations for elongated lanes, reducing detection accuracy. LaneATT[19] addresses this issue by noting the high correlation between lane positions and introducing global features to infer lane position, especially in cases of occlusion or missing lane markers. LaneATT[19] introduced an anchor-based attention mechanism that significantly improves detection speed and accuracy by utilizing global information. CLRNet[20] demonstrated the complementary roles of low- and high-dimensional features in lane detection, proposing Cross Layer Refinement (CLR) to dynamically adjust anchor points through network learning, removing the limitations of fixed anchors to optimize lane detection further. FLAMNet[21] proposes an adaptive line anchor framework integrated with a dual-branch network design to improve the accuracy and adaptability of capturing intricate lane geometries. By leveraging the Decomposed Self-Attention (DSA) mechanism alongside a Horizontal Feature Fusion Module (HFFM), the model enhances the perceptual range of line anchors in lane detection tasks. This approach strengthens robustness against occlusions and incomplete lane markings in challenging environments.
@@
Another approach employs domain adaptation techniques to address the style gap between foggy scenes and clear-weather images[24, 25]. Hnewa et al.[26] assumed that there is domain shift between images captured in normal and adverse weather conditions, proposing a multi-scale domain-adaptive YOLO that supports domain adaptation at different levels in the feature extraction stage. Li et al.[27] proposed a novel transfer learning-based framework for fog-adaptive object detection, which integrates dual-level adaptation (image-level and object-level) to mitigate domain shifts caused by foggy conditions. Notably, the method only requires labeled clear-weather images and unlabeled foggy data for training, eliminating the dependency on annotated foggy datasets.
@@ 2 Related Work
2.1 Segmentation-based Lane Detection
@@
2.2 Curve-based Lane Detection
@@
2.3 Anchor-based Lane Detection
@@
2.4 Row-wise-based Lane Detection
@@
3 Data Constrution
Due to the data-driven nature of deep learning models, the quality, size, and diversity of the datasets largely determine the applicability and performance of such models[36]. In the task of lane detection, models trained solely on clear weather data struggle to adapt to low-visibility foggy conditions. Issues such as blurred lane boundaries, faded lane colors, and reduced image quality often arise in fog, making lane detection more challenging. Therefore, training with a foggy weather lane dataset can significantly improve an autonomous driving system’s ability to handle such conditions, enhancing both detection accuracy and overall reliability[37, 38]. However, there is currently no publicly available foggy weather lane detection dataset. To address this gap, we have collected and constructed a foggy weather lane detection dataset called FoggyLane. In addition, we generated two supplementary datasets, FoggyCULane and FoggyTusimple, by using fog modeling techniques on existing CULane [12] and Tusimple [28] datasets to test the generalizability of our methods.
@@ 3 Data Constrution
3.1 FoggyLane
In the field of intelligent connected vehicle ADAS systems, existing lane detection algorithms have shown good performance under favorable weather conditions. However, their performance remains significantly inadequate under adverse weather conditions, such as fog. Given that fog is a high-risk scenario for traffic accidents, the limited visibility drastically reduces the driver’s reaction time and decision-making ability, making accidents more likely. However, current benchmark datasets for lane detection, including CULane[12], Tusimple[28], and LLAMAS[40], do not include data from foggy scenarios. To address this gap, we have developed a lane detection dataset, FoggyLane, specifically designed for foggy conditions. Given the randomness of foggy weather, data collection was carried out using a monocular forward-facing camera at different times and locations. This includes the morning of October 26, 2023, in Suzhou, Jiangsu Province, China, and two sessions in Urumqi, Xinjiang Uygur Autonomous Region, China: the morning of December 9, 2023, and the night of January 2, 2024. In these selected representative traffic scenarios, we sampled one frame every frames, resulting in a total of 1,086 RGB images: 444 images (resolution: $1920\times 1080$), 102 images (resolution: $2560\times 1920$), and 540 images (resolution: $2304\times 1296$). To simplify operations and enhance usability during training, we resized all images to a uniform resolution of $1640\times 590$.
However, since the duration of foggy weather conditions is limited, the number of collected foggy weather images is restricted. To improve the model’s performance, we extended our dataset by incorporating additional foggy driving footage. Specifically, we downloaded two foggy driving videos from Bilibili[41], extracting road-driving segments at 20-frame intervals, obtaining 44 and 293 frames, respectively. This added a total of 337 additional RGB images to the dataset. The original resolution of these images was $1920\times 1080$, and they were resized to $1640\times 590$ for consistency.
@@ 3.1 FoggyLane
Meanwhile, We compared FoggyLane with existing lane detection datasets. As shown in Table 1, the Caltech-Lanes and Tusimple datasets only cover single traffic scenarios. The LLAMAS [40] dataset focuses on highway scenarios, but its annotations are generated from high-precision maps rather than manually annotated. While the ApolloScape [42] and BDD100k [43] datasets contain multiple traffic scenarios, they lack instance-level annotations, and ApolloScape [42] does not provide publicly available test set annotations. In comparison, although CULane [12] and CurveLanes [44] include a wide range of traffic scenes with comprehensive annotations, they still lack foggy weather scenarios. For lane detection tasks, models trained only on clear weather conditions struggle to handle the challenges posed by fog, such as blurred lane boundaries, faded colors, and reduced image quality. To address these limitations, we developed FoggyLane, a comprehensive foggy weather lane detection dataset that includes 1,423 images featuring various challenging scenarios such as foggy roads during the day, curves, crossroads, and night fog. The dataset is meticulously annotated using the LabelMe tool, providing instance-level annotations, and it is available in Tusimple [28] and CULane [12] formats for broader accessibility.
Table 1: Comparison Between Our Dataset and Existing Lane Detection Datasets
@@ 3.2 FoggyCULane and FoggyTusimple
4 Methods
4.1 Overall structure
@@
4.2 Global Feature Fusion Module
In recent years, Transformer models have achieved significant success in natural language processing, leading researchers to explore their applications in computer vision. With self-attention mechanisms, Transformer models can capture dependencies between different regions and pixels within an image, thus facilitating a better understanding of global image information[52]. As a result, Transformer models have demonstrated outstanding modeling capabilities and performance improvements in various traffic vision tasks[53, 54, 55, 56]. However, applying Transformers to the visual domain still faces some challenges. Firstly, traditional Transformer architectures lack the hierarchical structure and down-sampling operations found in CNN models, making them less adaptable to changes in resolution and receptive field, which may hinder handling of complex visual entities. Additionally, calculating global attention at high resolutions can be computationally intensive.
@@ 4.2 Global Feature Fusion Module
4.3 Kernel Feature Fusion Module
@@
4.4 Low-level Edge Enhanced Module
Images in foggy conditions differ from those captured under good lighting primarily due to their blurriness and low contrast. This effect renders the edges of objects and details within the scene indistinct and hard to detect, impacting the algorithm’s ability to understand and recognize the scene accurately. Edge information, beyond simply marking object boundaries, also implicitly contains structural and spatial information about objects, forming a critical foundation for tasks such as object detection, segmentation, and tracking. Using the Canny edge detection algorithm with identical threshold settings, we processed both foggy and normal images, as shown in the Fig. 10. It is evident that fog significantly degrades edge information, particularly affecting lane line edges, which become nearly indiscernible. Given that lane detection relies heavily on edge information—lane positioning is typically determined by the boundary between lane edges and the road background—both global scene information and detailed edge information are essential for accurate lane detection.
@@ 4.4 Low-level Edge Enhanced Module
Table 2: Comparison with Advanced Lane Detection on FoggyLane dataset.
@@
Specifically, we add an auxiliary path in the network to learn the edge features of the input image. First, we apply the Canny operator to the input image for edge detection, supplemented by lane line labels to reconstruct any missing lane edges, generating the training labels for the edge enhancement module. To avoid interfering with high-level semantic extraction and subsequent detection segmentation, while strengthening the low-level network’s learning of edge textures and other low-dimensional features, we introduce a branch at the second stage of the backbone network. This branch sequentially applies a $3\times 3$ and a $1\times 1$ convolution. We then compute the cross-entropy loss between the branch output and the edge labels. The detailed structure is illustrated in the bottom-left of the Fig. 8. Notably, this auxiliary enhancement path is only active during training, thus not increasing computation or processing time during deployment.
@@
Table 3: Comparison with Advanced Lane Detection on FoggyCULane dataset
@@ 5.1 Datasets
FoggyLane is a lane detection dataset specifically designed for foggy scenarios, containing 1,423 frames of foggy lane images for training and testing. All images are of resolution $1640\times 590$ pixels.
FoggyCULane is synthesized from the large-scale challenging lane detection dataset CULane using a foggy image modeling approach. It maintains the same number of images as CULane, which includes 88,880 images for training and 34,680 images for testing. The resolution of all images is also $1640\times 590$ pixels.
FoggyTusimple is synthesized from one of the most widely used benchmark datasets for lane detection, Tusimple, through the foggy image modeling method. At the same time, the dataset follows the same structure as Tusimple, comprising 3,268 training images, 358 validation samples, and 2,782 test cases. All visual data maintains a consistent resolution of $1280\times 720$ pixels.
Figure 13: Several metrics of our model’s performance during the training process on the FoggyLane and FoggyTusimple datasets, with results on FoggyLane on the left and FoggyTusimple on the right.
5.2 Evaluation Metric